# Concept Based Explanations and Class Contrasting

### v1

@@ 5 Class Contrasting
In this section we look at contrasting classes, i.e. investigating why the model predicts class ’A’ over ’B’. A problem that can arise from only explaining each class individually, is that the explanation for two classes can be very similar (e.g. Alaskan Malamute and Siberian Husky, see Section 6), where the explanations look virtually indistinguishable. In that example, the explanation for Alaskan Malamute actually fails (for the dataset visualization the model predicts an average of 0.14 for Alaskan Malamute, and the majority class is Eskimo Dog with a prediction of 0.37, and the prediction for Siberian Husky is 0.16). So there we rather get an explanation for a more generic dog. Providing an explanation (six prototypes again) that contrast Alaskan Malamute with Siberian Husky, provides a better explanation for Alaskan Malamute compared to looking at that class in isolation. For that new explanation the model predicts 0.50 for Alaskan Malamute.
@@
Model Default Pred Shifted Pred Pred Swap 0.1 Pred Swap 0.2 Pred Swap 0.5
ResNet34 Digipath 0.018 ($\pm$0.021) 0.458 ($\pm$0.342) 69.4% 65.2% 49.0%
ResNet50 Robust 0.001 $\pm$ 0.000 ($\pm$0.004) 0.637 $\pm$ 0.004 ($\pm$0.392) 79.5% $\pm$ 0.2 73.9% $\pm$ 0.3 60.4% $\pm$ 0.4
ResNet34 $0$$\pm$ 0.000 ($\pm$0.003) 0.263 $\pm$ 0.004 ($\pm$0.382) 29.1% $\pm$ 0.3 27.4% $\pm$ 0.3 23.1% $\pm$ 0.4
ResNet50 $0$$\pm$ 0.000 ($\pm$0.002) 0.335 $\pm$ 0.005 ($\pm$0.411) 34.3% $\pm$ 0.6 32.8% $\pm$ 0.5 29.5% $\pm$ 0.6
@@ 5.1 Method
The idea is to train a linear classifier (single linear layer with a scalar bias and a sigmoid in the end) for binary classification between activations that are predictive for class ’A’ and those that predictive for class ’B’. Then the normal of the hyperplane spanned by the linear classifier would ideally point from class ’A’ to ’B’, and those activations that are far along the hyperplane normal in the direction of class ’A’ would be predictive for class ’A’ vs ’B’. Its a similar approach as before for explaining a single class, the only difference is that instead of using attribution to filter and score the activations, we use the linear classifier to filter and score the activations. The rest is the exact same. Figure 3 shows our approach to generate the hyperplane separating class ’A’ and ’B’. This hyperplane is used in the next step shown in Figure 4 to extract the n most important concepts separating class ’A’ from ’B’.
@@ 5.1.1 Hyperplane
First we describe Figure 3, which shows the generation process of the linear classifier. Step 1 to 3 are done individually per class, step 4 is run for all combinations of any two classes. In the case of ImageNet with 1000 classes we would get roughly one million linear classifier (we exclude the combination of a class with itself). Step 1 and 2 are similar to Figure 2, the only difference is that we do not scale the activations by the attribution, but filter only. And the filtering is a bit more aggressive. Instead of filtering out only those pixels that have negative attribution we also filter out all those pixels that have less than 0.25 of the maximum attribution over all pixels of that image. The reason why we choose a higher cutoff, is that previously the activations would be scaled by the attribution, so that those with lower attribution would have lower impact on the final prototypes. But now all the activations that pass the filtering step have the same impact on the linear classifier, therefore we use a higher cutoff.
In the next step the activations of the pixels remaining after the filtering are saved (i.e. we save a list of vectors per class, each vector representing a pixel in the activations at layer Y). Then in the last step we train a linear classifier to do binary classification between two lists of vectors, that were created in step three (one list for class ’A’ and one for class ’B’). The linear classifier is always trained for 100 epochs, running the whole dataset at once (no mini-batching). As optimizer we use stochastic gradient descent (Robbins & Monro, 1951), (Bottou, 2010) with a learning rate of 0.01.
@@ 5.1.2 Extract Separating Concepts
Now we describe how we extract the six prototypes for the contrasting. Figure 4 shows our method to find those six concepts separating class ’A’ from ’B’. The method is similar to the method explaining a single class, shown in Figure 2. The only difference is in step 2, where we now scale and filter the activations by the linear classifier, instead of using attribution.
We run the linear classifier (without the sigmoid, so only the dot product between the weight vector and the pixel and adding the scalar bias) pixel wise on all the intermediate activations we get at layer Y. If $w$ is the weight vector, $b$ is the scalar bias and $x$ is a pixel of the intermediate activations at layer Y, then we do: $z=w\cdot x+b$, where the output $z$ is a scalar value. Then for each pixel we do: $x^{\prime}=\text{max}(z,0)\cdot x$, that is we scale the activations by $z$ if $z$ is larger than zero, otherwise we set those activations to zero. We only keep those activations that are on the side of the hyperplane that is predictive for class ’A’, and scale those remaining activations by how far along the hyperplane normal they are. The remaining two steps are the exact same as for explaining a single class. The modified activations are decomposed via NMF into six concepts, and those six concepts are then visualized.
Figure 5: Visualization for the 6 NMF components for Alaskan Malamute from dataset crops
Figure 6: Visualization for the 6 NMF components for Siberian Husky from dataset crops
Figure 7: Visualization for the 6 NMF components pro Alaskan Malamute vs Siberian Husky from dataset crops
Figure 8: Visualization for the 6 NMF components pro Siberian Husky vs Alaskan Malamute from dataset crops
@@ 5.2.1 Shifting Test
The test we propose for the contrasting, tests the direction of the hyperplane. Ideally, the hyperplane normal should point from class ’B’ towards class ’A’. This means that if we would start from an image of class ’B’, and then translate the activations in the hidden layer along the hyperplane normal, we would like the model to swap prediction to class ’A’. We do the translation using ten offset values (those are fixed and the same for all classes and models) and report the result of the offset value where the prediction for class ’A’ is highest.
The results from this translation test are shown in Table 3. In the following we describe how to read that table. Default pred as well as shifted pred both show the results for class ’A’. Default pred shows the mean prediction for class ’A’ for the original activations (when not adjusting anything). Shifted pred shows the mean prediction for class ’A’ when translating the activations along the hyperplane normal. For the digipath model we run all combination of classes (there we only have 37 classes), and report the mean ($\pm$ standard deviation over all samples). For the imagenet models we have too many classes to run all combinations time wise, instead there we use each of the 1000 classes as starting class (’B’), and for each of them we choose only 10 random classes (’A’) to shift towards. Then we repeat this 5 times and report mean $\pm$ standard deviation over the 5 runs as well as ($\pm$ standard deviation over all samples).
The next three columns show in how many of the class combinations the model swaps prediction (i.e. for those shifted activations the model now predicts class ’A’ as majority class instead of class ’B’). Also the model needs to predict class ’A’ with at least a certain average confidence. For example, for pred swap 0.5 it would need to predict at least 0.5 for class ’A’, on average over all samples.
We can see that for all the models, the original prediction for class ’A’ is very low (default pred column). Which makes sense, since we choose images where the model is predicting another class ’B’ as majority class. But the shifted prediction becomes much larger. When comparing the shifted prediction with the default prediction, we have a minimum increase factor of 25x for the ResNet34 Digipath model, and a maximum increase factor of 1494x for the ResNet50 model. We get a minimum average shifted prediction of 0.263 for the ResNet34 model, and a maximum of 0.637 for the ResNet50 Robust model.
Table 4: Prediction for the original and stitched images. Both when inserting the person, as well as inserting zeros instead of the person.
Class Original Pred Person Inserted Pred Black Inserted Pred
Siberian Husky 0.377 0.265 0.342
Alaskan Malamute 0.203 0.265 0.189
@@ 6.1 Alaskan Malamute vs Siberian Husky
In this section we investigate why the ResNet50 Robust model classifies an image as Alaskan Malamute vs Siberian Husky and vice versa (in the appendix we show more examples also from the other models). We apply the method at layer3.5. Both classes are dog breeds that look very similar. Figure 5 shows visualizations from dataset crops for those six most important concepts supporting Alaskan Malamute, whereas Figure 6 shows the same for Siberian Husky. We can see that the model mostly uses the dog nose, eyes and ears to predict either class. For example, the first concept for Siberian Husky shows dog noses, the second looks unclear (from the dataset visualization it looks like fur), the third shows eyes, the fourth looks like it shows fur at the head, the fifth shows the top of head without ears and the last shows pointy ears. But both sets of images (Alaskan Malamute vs Siberian Husky) look very similar, just based on this it is not possible to understand why the model predicts one class versus the other.
@@
Next we can contrast both classes (using our method described in Section 5), and look at the six most important concepts separating one class from the other. Now we can see differences between the visualizations of both classes. For Alaskan Malamute, there is now only one component showing part of the dog head (the full head without ears in this case), as shown in the second row of Figure 7. For Siberian Husky, most of the components still show parts of the dog head, the bottom most component looks a bit different though. Now it shows larger eyes. For the Alaskan Malamute, the three bottom most components are interesting (since they were also not present when looking only at the Alaskan Malamute class in isolation), shown in Figure 7. The top of those bottom three components shows dog feet on grass, the next one shows grass and the last one looks like the top of human heads. The last two components (grass and top of human heads), point to a bias in the model, cause whether the dog is standing on grass or snow, or whether a human is next to the dog or not, does not change the class the dog belongs to. So the model should not rely on those features to separate between the two dog classes.
@@
To test this bias, we insert an image of a person into each of the sixty images predicted as Siberian Husky (always at the bottom right), and then check whether the model changes its prediction to Alaskan Malamute. Something to note is that the person is cropped out of a (validation) image where the model predicts Alaskan Malamute. One of those sixty modified images is shown in Figure 9, once with the person inserted and once with zeros inserted. The image is shown after normalization.
@@
Figure 9: Example of inserting the person (or zeros) into an originally as husky predicted image
@@
Table 4 shows the average prediction for Siberian Husky and Alaskan Malamute for those sixty images. Originally the model predicts more Siberian Husky (0.377). After inserting the person into the lower right of each of the images, the model predicts the same amount of Siberian Husky and Alaskan Malamute (both 0.265). Since we occlude part of the image by inserting the person, we also test the prediction if we insert zeros (black) instead of the person. In that case the prediction does not change much, and slightly decreases for both Siberian Husky as well as Alaskan Malamute.
7 Additional Discussion
The method to explain a single class could be seen as an interpretability method (which was our goal), but can also been seen as a method to generate high level adversarial examples. That is, because it provides an automated method to extract crops from six images where the model is not predicting the targeted class, then for the combined image the model is often predicting that targeted class. This test (or attack) works worse for the ResNet50 Robust model which has adversarially robust weights (it was trained for robustness against gradient descent attacks). While it still works for 42.9% of the classes, the confidence is lower (the average prediction for the targeted class is only 0.190). Also, when it predicts the targeted class, then only for 11.3% of the classes does it do so with a confidence of larger than 0.5. As comparison, the ResNet50 and ResNet34 model with weights from pytorch model zoo do so for 62.6% respectively 55.4% of the classes (much more often). For the ResNet50 model with weights from pytorch model zoo the test (or attack) works for 71.0% of the classes and the confidence for the targeted class is higher, there we have an average prediction of 0.619. Similar for the ResNet34 model with weights from pytorch model zoo, there the test works for 65.4% of the classes with an average prediction of 0.547.
While using the last convolutional layer (layer4.2) provides best results in the test, the explanations may be less interpretable compared to using layer3.5.
@@ 8 Limitations
We only tested the method on convolutional neural networks. Further, for a class where the test fails, the method should likely not be trusted. Also we need to manually choose a hidden layer in the model where the method should be applied at (we could potentially run the prediction test for the prototypes for different layers of the model, and then choose the layer where the test performs best).
@@ 9 Acknowledgments
R.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projectnumber 459360854 (DFG FOR 5347 Lifespan AI). D.O.B. acknowledges the financial support by the Federal Ministry of Education and Research (BMBF) within the T!Raum-Initiative ”#MOIN! Modellregion Industriemathematik” in the sub-project ”MUKIDerm”.
@@ Appendix A Hardware and Software
On the software side we use pytorch (Paszke et al., 2019) version 1.13.1 with cuda version 11.6 and torchvision (TorchVision maintainers and contributors, 2016) version 0.14.1. For the attribution methods (for DeepLift and SmoothGrad) we use the captum library (Kokhlikyan et al., 2020) version 0.6.0. For NMF we use scipy (Virtanen et al., 2020) version 1.10.1.
@@
Appendix B Generate Visualizations via Gradient Descent

### v2

@@ 5 Class Contrasting
In this section we look at contrasting classes, i.e. investigating why the model predicts class ’A’ over ’B’. A problem that can arise from only explaining each class individually, is that the explanation for two classes can be very similar (e.g. Alaskan Malamute and Siberian Husky, see Section 6), where the explanations look virtually indistinguishable. In that example, the explanation for Alaskan Malamute actually fails (for the dataset visualization the model predicts an average of 0.26 for Alaskan Malamute, and the majority class is Siberian Husky with a prediction of 0.42). Which means we only get an explanation for Siberian Husky, but none for Alaskan Malamute.
Therefore in the next step we do not explain Alaskan Malamute in isolation, but contrast it with the class it failed with (Siberian Husky). That results in a much better explanation for Alaskan Malamute, in the sense that the model now predicts it as majority class with a confidence of 0.81.
In the following we describe our method and results for such contrasting between two classes.
@@
Model Layer Default Pred Shifted Pred
ResNet50 4.2 $0$$\pm$ 0.000 ($\pm$0.002) 0.995 $\pm$ 0.000 ($\pm$0.036)
ResNet50 3.5 $0$$\pm$ 0.000 ($\pm$0.002) 0.375 $\pm$ 0.004 ($\pm$0.420)
@@ 5.1 Method
The idea is to train a linear classifier (single linear layer with a scalar bias and a sigmoid in the end) for binary classification between activations that are predictive for class ’A’ and those that predictive for class ’B’. Then the normal of the hyperplane spanned by the linear classifier would ideally point from class ’A’ to ’B’, and those activations that are far along the hyperplane normal in the direction of class ’A’ would be predictive for class ’A’ vs ’B’. Its a similar approach as before for explaining a single class, the only difference is that instead of using attribution to filter and score the activations, we use the linear classifier to filter and score the activations. The rest is the exact same. Figure 3 shows our approach to generate the hyperplane separating class ’A’ and ’B’. This hyperplane is used in the next step to extract the n most important concepts separating class ’A’ from ’B’.
@@ 5.1.1 Hyperplane
First we describe Figure 3, which shows the generation process of the linear classifier. Step 1 to 3 are done individually per class, step 4 is run for all combinations of any two classes. In the case of ImageNet with 1000 classes we would get roughly one million linear classifiers (we exclude the combination of a class with itself). Step 1 and 2 are similar to Figure 2, the only difference is that we do not scale the activations by the attribution, but filter only. And the filtering is a bit more aggressive. Instead of filtering out only those pixels that have negative attribution, we also filter out all those pixels that have less than 0.25 of the maximum attribution over all pixels of that image. The reason why we choose a higher cutoff, is that previously the activations would be scaled by the attribution, so that those with lower attribution would have lower impact on the final concepts. But now all the activations that pass the filtering step have the same impact on the linear classifier, therefore we use a higher cutoff.
In the next step the activations of the pixels remaining after the filtering are saved (i.e. we save a list of vectors per class, each vector representing a pixel in the activations at layer Y). Then in the last step we train a linear classifier to do binary classification between two such lists of vectors (one list for class ’A’ and one for class ’B’). The linear classifier is always trained for 100 epochs, running the whole dataset at once (no mini-batching). As optimizer we use stochastic gradient descent Robbins and Monro (1951), Bottou (2010) with a learning rate of 0.01.
Figure 5: Visualization for the 4 NMF components for Alaskan Malamute
Figure 6: Visualization for the 4 NMF components pro Alaskan Malamute vs Siberian Husky
Figure 7: Visualization for the 4 NMF components for Siberian Husky
Figure 8: Visualization for the 4 NMF components pro Siberian Husky vs Alaskan Malamute
@@ 5.1.2 Extract Separating Concepts
Now we describe how we extract the n concepts for the contrasting. The method is similar to the method explaining a single class, shown in Figure 2. The only difference is in step 2, where we now scale and filter the activations by the linear classifier, instead of using attribution.
We run the linear classifier (without the sigmoid, so only the dot product between the weight vector and the pixel and adding the scalar bias) pixel wise on all the intermediate activations we get at layer Y. Let $w$ be the weight vector, $b$ the scalar bias and $x$ the pixel of the intermediate activations at layer Y, then we do: $z=w\cdot x+b$, where the output $z$ is a scalar value. Then for each pixel we do: $x^{\prime}=\text{max}(z,0)\cdot x$, that is we scale the activations by $z$ if $z$ is larger than zero, otherwise we set those activations to zero. We only keep those activations that are on the side of the hyperplane that is predictive for class ’A’, and scale those remaining activations by how far along the hyperplane normal they are. The remaining two steps are the exact same as for explaining a single class. The modified activations are decomposed via NMF into n concepts, and those n concepts are then visualized.
@@ 5.2.1 Shifting Test
The test we propose for the contrasting, tests the direction of the hyperplane. Ideally, the hyperplane normal should point from class ’B’ towards class ’A’. This means that if we would start from an image of class ’B’, and then translate the activations in the hidden layer along the hyperplane normal, we would like the model to swap prediction to class ’A’. We do the translation using ten offset values (those are fixed and the same for all classes and layers) and report the result of the offset value where the prediction for class ’A’ is highest. We do not simply use a single large value, because the prediction for class ’A’ may decrease again when going too far along the hyperplane normal (this happens for layer3.5).
The results from this translation test are shown in Table 3. Default pred as well as shifted pred both show the results for class ’A’. Default pred shows the mean prediction for class ’A’ for the original activations (when not adjusting anything). Shifted pred shows the mean prediction for class ’A’ when translating the activations along the hyperplane normal. We have too many classes to run all combinations time wise, instead we use each of the 1000 classes as starting class (’B’), and for each of them we choose only 10 random classes (’A’) to shift towards. Then we repeat this 5 times and report mean $\pm$ standard deviation over the 5 runs as well as ($\pm$ standard deviation over all samples).
Table 4: Prediction for the original and stitched images. Both when inserting the grass patch, as well as inserting zeros instead of that grass patch
Class Original Pred Grass Inserted Black Inserted
Siberian Husky 0.165 0.117 0.143
Alaskan Malamute 0.080 0.100 0.071
concept 0 concept 1 concept 2 concept 3
Siberian Husky 0.176 0.162 0.444 0.117
Alaskan Malamute 0.127 0.155 0.325 0.080
@@ 6.1 Alaskan Malamute vs Siberian Husky
In this section we investigate why the ResNet50 Robust model classifies an image as Alaskan Malamute vs Siberian Husky and vice versa (in the appendix we show more examples). We apply the method at layer4.2. Both classes are dog breeds that look very similar. Figure 6 shows visualizations from dataset crops for the four most important concepts supporting Alaskan Malamute, whereas Figure 8 shows the same for Siberian Husky. We can see that the model mostly uses the dog face to predict either class. The explanation for Alaskan Malamute is a failure case, since the model predicts Siberian Husky as majority class for the four concepts (0.26 for Alaskan Malamute and 0.42 for Siberian Husky). Both sets of images (Alaskan Malamute and Siberian Husky) look very similar, just based on this it is not possible to understand why the model predicts one class versus the other.
@@
Next we can contrast both classes (using our method described in Section 5), and look at the four most important concepts separating one class from the other. Now we can see differences between the visualizations of both classes, and the contrasted explanation (Figure 6) gets predicted as Alaskan Malamute with a confidence of 0.81 (much higher compared to the 0.26 when explaining Alaskan Malamute in isolation). For Alaskan Malamute, there is now only one component showing the dog head, as shown in the third row of Figure 6. Interestingly, three out of four components show grass, which potentially points to a bias in the model, since it should not matter whether the dog is standing on grass or e.g. snow.
@@
To test this bias, we insert an image of grass into the first 128 training images predicted as Siberian Husky (always at the bottom right), and then check whether the model changes its prediction to Alaskan Malamute. One of those 128 modified images is shown in Figure 9, once with the grass image inserted and once with zeros inserted. The image is shown after normalization.
@@
Figure 9: Example of inserting the grass patch (or zeros) into an originally as husky predicted image (image shown after normalization)
@@
Table 4 shows the average prediction for Siberian Husky and Alaskan Malamute for those 128 images. Originally the model predicts more Siberian Husky (0.165). After inserting the grass image into the lower right of each of the images, the model predicts less Husky (0.117) and more Malamute (0.100 vs originally 0.080). Since we occlude part of the image by inserting the grass image, we also test the prediction if we insert zeros (black) instead. In that case the prediction does not change much, and slightly decreases for both Husky as well as Malamute.
The bottom half of the table reports the prediction when inserting a concept from contrasting Malamute with Husky (from Figure 6). E.g. concept 0 would report the prediction when inserting the first image of the first row from Figure 6, concept 1 would be the first image from the second row. concept 1 (the tail image) works best, inserting it slightly decreases the prediction for Husky and raises the prediction for Malamute close to that of Husky.
@@ 7 Limitations
We only tested the method on convolutional neural networks. Further, for a class where the test fails, the method should not be trusted (as shown in the case study where it fails for Malamute). That test fails in 9% of the classes for the worst model. The stricter test (excluding dataset crops predicted as the target class) fails in 9% of the classes for the best model, and for the worst model it fails in 14% of the classes. That stricter test also requires a dataset that contains a variety of concepts from different classes. Because it will likely not work for e.g. binary classification between two very dissimilar classes (like airplane vs dog). Because its likely not possible to visualize concepts that explain the prediction for dog from airplane images.
Also, visualizing the concepts via dataset examples (as opposed to using iterative gradient descent to synthesize them) may result in concepts that are misleading. For example, if a concept shows a dog head, does it mean that concept represents the whole head, or maybe only the nose or eyes? We discuss that more in Appendix E, F.
8 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning, especially in the field of explainable AI. Obtaining a better understanding of the behavior of the models could have a positive effect.
@@ 9 Acknowledgments
R.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projectnumber 459360854 (DFG FOR 5347 Lifespan AI). D.O.B. acknowledges the financial support by the Federal Ministry of Education and Research (BMBF) within the T!Raum-Initiative "#MOIN! Modellregion Industriemathematik" in the sub-project "MUKIDerm".
@@ Appendix A Hardware and Software
On the software side we use pytorch Paszke et al. [2019] version 1.13.1 with cuda version 11.6 and torchvision TorchVision maintainers and contributors [2016] version 0.14.1. For the attribution methods (for DeepLift, SmoothGrad and gradient times activation) we use the captum library Kokhlikyan et al. [2020] version 0.6.0. For NMF we use scipy Virtanen et al. [2020] version 1.10.1.
@@
Appendix B Computational Requirements
In the following, all times are reported running on a single RTX 2080Ti GPU with an AMD EPYC 7262 8-Core Processor (32 CPUs). Extracting the NMF components for all 1000 classes for a single model and layer takes 15 hours for the ResNet50 model. Visualizing the concepts for all 1000 classes adds another 14 hours (around 50 seconds per class), here the performance cost comes from comparing each of the n NMF components with each of the 11.5 million dataset crops (vector wise cosine similarity).
For the class contrasting, generating the activations for all 1000 classes takes 3 hours (step 1 to 3 shown in Figure 3). Training the pairwise linear classifiers takes 35 hours. Using those trained classifiers to contrast two classes then takes around 90 seconds (40 seconds for generating the NMF components and another 50 seconds for the visualizing the concepts).
Appendix C concept Intruder Detection
Table 5: Detecting an intruder concept
Number of Components Detection Accuracy
4 77%
8 76%
Figure 10: Intruder detection experiment
In order of testing the human interpretability of the extracted concepts, we asked 4 persons affiliated with the authors of this paper (co-workers) to detect an intruder image out of five images. They are shown five images, from which four represent one concept and one represents another concept but from the same class. The goal is to identify that one image that represents the other concept. The idea is to check how human interpretable the concepts are, i.e. if they all look the same, then their usefulness as explanation will be low. This is done for fifty intruder images (fifty steps), each time selecting a new random class out of the 1000 classes and two random concepts (one for the four images and another concept for the intruder image). And this experiment is done using the ResNet50 model and layer4.2.
The result is shown in Table 5. Its run twice, one time using an explanation with four concepts and one time using an explanation with eight concepts. This was to check whether the explanation becomes less interpretable if more concepts are used (eight vs four). Because there may be more overlap between eight concepts than there is with four. But the result is similar (77% detection accuracy for four concepts and 76% for eight concepts). Random guessing would be around 20%, therefore we see that as a confirmation that our method provides human interpretable explanations.
Appendix D Hyperparameters
D.1 Number of Samples
Figure 11: Testing different number of components in the NMF and different hidden layers
In Figure 11 we plot how our method performs depending on the number of samples that is used to generate the explanation. We vary the number of samples between 50, 100 and then in steps of 100 till 900. We keep the other hyperparameters (layer4.2, 4 components, DeepLift as attribution method) fixed. The left plot shows the average prediction for the class that should be explained and the right plot shows how often the explanation is predicted as that target class as majority class. And we exclude patches that are predicted as the target class in sampling the closest dataset patches for the explanation.
The standard deviation in the prediction for 500 samples is 0.02 (as shown in Table 2). Assuming its the same for the other data points as well, then the results for 100 to 900 samples are roughly the same, with the result for 50 samples being a bit lower. Therefore, we would recommend to have at least 100 samples available per class.
D.2 Attribution Method
Table 6: Prediction test for the explanation. Allowing dataset crops predicted as the target class
Model Layer Attribution Method Samples Average Pred Predicts Target Class
ResNet50 4.2 DeepLift 50 0.976 98.0%
ResNet50 4.2 DeepLift plus SmoothGrad 50 0.972 99.0%
ResNet50 4.2 Gradient times Activation 50 0.975 98.0%
ResNet50 3.5 DeepLift 500 0.524 56.0%
ResNet50 3.5 Gradient times Activation 500 0.497 57.0%
Table 7: Prediction test for the explanation. Excluding dataset crops predicted as the target class
Model Layer Attribution Method Samples Average Pred Predicts Target Class
ResNet50 4.2 DeepLift 50 0.860 88.0%
ResNet50 4.2 DeepLift plus SmoothGrad 50 0.848 87.0%
ResNet50 4.2 Gradient times Activation 50 0.860 88.0%
ResNet50 3.5 DeepLift 500 0.837 85.0%
ResNet50 3.5 Gradient times Activation 500 0.799 85.0%
In Table 6 and Table 7 we investigate the effect of using different attribution methods. This experiment is run using the ResNet50 model for layer4.2 and layer3.5. As attribution methods we use DeepLift, DeepLift plus SmoothGrad (using 20 samples and a standard deviation in the gaussian noise of 0.25) and gradient times activation. In layer4.2 all three methods perform similar, with DeepLift plus SmoothGrad performing slightly worse. In layer3.5 we only compare DeepLift and gradient times activation, here we want to check whether the non-linearity becomes a problem when using gradient times activation. But gradient times activation only performs slightly worse compared to using DeepLift (slightly lower prediction). Overall, simply using gradient times activation can also be used, especially in the penultimate layer (layer4.2), where there is no non-linearity anyway.
D.3 Shifting Test Prediction over Offset
Figure 12: Plotting the average prediction for class ’A’ (the target class). The x-axis shows how far the activations are shifted along the hyperplane normal pointing from class ’B’ towards ’A’
Figure 12 illustrates the offsets used in the shifting test in Table 3 in the main paper. For layer4.2, we could simply use a very large offset value, but for layer3.5 the prediction decreases again when the offset value is becomes too large.
D.4 Number of Components and Hidden Layer
Figure 13: Testing different number of components in the NMF and different hidden layers. The target class is allowed in sampling of the concepts
Figure 13 shows the same results as 4 in the main paper, but allows concepts that are predicted as the target class. The result is similar, the deeper the layer the better and our method slightly outperforms CRAFT. And again layer1.2 and layer2.3 do not work (the prediction is very low).
Appendix E Generate Visualizations via Gradient Descent
For this section and the next section F we use an in-house model trained to segment tumor between 37 classes in digital pathology tissue images. It uses a modified UNet with a ResNet34 backbone and in the following we use layer3.2 (that would be 21 layers into the ResNet34). The dataset images have a size of 1536x1536 and the concepts have a size of 512x512. We use 60 samples (instead of 500) and as attribution method we use DeepLift plus SmoothGrad. And we use $n=6$ components in the NMF (i.e. we use six concepts in the explanation).